A builder claims that bidets are installed in more than 50% of all homes being constructed today in suburban Lapu-lapu. Would you agree with this claim if a random survey of new homes in this city showed that 54 out of 150 had bidets installed? Use a 0.05 level of significance.


1
Expert's answer
2021-10-03T17:29:48-0400

Solution:

Null hypothesis, H0 : "p=0.50"

Alternate hypothesis, H1 : "p>0.50"

"\alpha=0.05", two -tailed test.

Critical value"=z_{\alpha/2}=1.96"

Test statistic"=z=\dfrac{\hat p-p}{\sqrt{pq/n}}\sim N(0,1)"

"\hat p=54/150=0.36"

Then, "z=\dfrac{0.36-0.50}{\sqrt{0.50\times0.50/150}}=-3.43"

Thus, we have enough evidence to reject the null hypothesis.

So, "p>0.50".